# Location–Scale Calibration for Generalized Posterior

### v2

@@ 1 Introduction
In this paper, we revisit the open-faced sandwich adjustment (shaby2014open-faced) and place it in a general asymptotic framework for generalized Bayes based on loss functions with an effective sample size, allowing sample-size-dependent priors, possibly non-i.i.d. observations, and convex penalties under model misspecification. First, we show that, under sample-size-dependent priors, the generalized posterior admits a normal limit in total variation and we establish a prior–penalty correspondence that identifies the target curvature and variability. Second, building on this justification, we improve the open-faced sandwich adjustment from covariance rescaling to a location–scale calibration that is implementable via plug-in sandwich estimators, requiring only posterior draws and empirical moments (no bootstrap or learning-rate tuning). Third, we prove that the calibrated draws converge to the target distribution for any learning rate, so that the asymptotic uncertainty quantification is invariant to the choice of learning rate. Our empirical illustration for a random-intercept mixed model with Huber loss and Gaussian prior with sample-size-dependent scale shows that the calibrated intervals track frequentist benchmarks across several orders of magnitude in the learning rate, whereas uncalibrated generalized Bayes posteriors exhibit substantial sensitivity.
@@ 2 Problem setup
Fix $p\in\mathbb{N}$ and let $\Theta\subset\mathbb{R}^{p}$ be an open parameter space. Write $\boldsymbol{\theta}\in\Theta$ for the parameter, and let $\|\cdot\|$ denote the Euclidean norm. Let $M_{n}:\Theta\to(-\infty,\infty]$ be an empirical criterion with effective scale $s_{n}\to\infty$; typical examples include i.i.d. additive losses ($s_{n}=n$), $m$-variate $U$-type losses ($s_{n}=\binom{n}{m}$ for non-degenerate kernels), and kernel-smoothed losses with bandwidth $h_{n}>0$ ($s_{n}=nh_{n}$, $h_{n}\to 0$ and $nh_{n}\to\infty$). Define the rescaled criterion and its first and second derivatives by
@@
We write $R(\boldsymbol{\theta}):=\mathbb{E}_{P^{\star}}[\bar{M}_{n}(\boldsymbol{\theta})]$ for the corresponding population criterion, where $P^{\star}$ denotes the true data-generating distribution.
@@
This loss-based representation makes explicit that the learning rate $\eta$ scales the entire loss, in line with bissiri2016general. Throughout the main text we assume that $\rho$ is $C^{2}$ in a neighborhood of the target point introduced below (see Assumption 2.1); the convex non-smooth case is treated in Section D.
Let $\{\mathcal{D}_{i}\}_{i=1}^{n}$ be observations from an unknown distribution $P^{\star}$ and $\Psi(\boldsymbol{\theta}):=\mathbb{E}_{P^{\star}}\!\left[\psi(\mathcal{D}_{1},\boldsymbol{\theta})\right]$ denote the population estimating function corresponding to $M_{n}$, for some measurable $\psi$. In the $M$-estimation case $M_{n}(\boldsymbol{\theta})=\sum_{i=1}^{n}m(\mathcal{D}_{i},\boldsymbol{\theta})$, one may take $\psi(\mathcal{D},\boldsymbol{\theta})=\nabla_{\boldsymbol{\theta}}m(\mathcal{D},\boldsymbol{\theta})$. The penalized population equation is
@@
$\boldsymbol{J}^{\star}:=\nabla_{\boldsymbol{\theta}}\Psi\bigl(\boldsymbol{\theta}^{\lambda}\bigr),\quad\boldsymbol{K}^{\star}:=\operatorname{Var}_{P^{\star}}\!\left[\psi\bigl(\mathcal{D}_{1},\boldsymbol{\theta}^{\lambda}\bigr)\right],\quad\boldsymbol{H}_{\rho}\bigl(\boldsymbol{\theta}^{\lambda}\bigr):=\nabla^{2}\rho\bigl(\boldsymbol{\theta}^{\lambda}\bigr).$
@@
There exists an open neighborhood $\mathfrak{N}$ of $\boldsymbol{\theta}^{\lambda}$ such that: (i) the penalized population equation (1) has the unique solution $\boldsymbol{\theta}^{\lambda}\in\mathfrak{N}$; (ii) $R\in C^{3}(\mathfrak{N})$ and $\sup_{\boldsymbol{\theta}\in\mathfrak{N}}\|\nabla^{3}\bar{M}_{n}(\boldsymbol{\theta})\|=O_{p}(1)$; (iii) $\sup_{\boldsymbol{\theta}\in\mathfrak{N}}\|\boldsymbol{J}_{n}(\boldsymbol{\theta})-\boldsymbol{J}^{\star}\|\to_{p}0$ and $\boldsymbol{J}^{\star}$ is nonsingular; and (iv) $\sqrt{s_{n}}\{\boldsymbol{U}_{n}(\boldsymbol{\theta}^{\lambda})-\Psi(\boldsymbol{\theta}^{\lambda})\}\to_{d}\mathcal{N}(\boldsymbol{0},\boldsymbol{K}^{\star})$.
@@
Let $\mathcal{U}_{n}=\{\boldsymbol{\theta}:\|\boldsymbol{\theta}-\boldsymbol{\theta}^{\lambda}\|\leq c/\sqrt{s_{n}}\}$ for some fixed constant $c>0$. The prior remainder $r_{n}$ is locally Lipschitz on $\mathcal{U}_{n}$ with Lipschitz constant $L_{n}=o_{p}(\sqrt{s_{n}})$.
@@
Under Assumptions 2.12.4, the following centers are admissible: (i) any maximum a posteriori (MAP) estimator, (ii) the posterior mean $\boldsymbol{\theta}_{\mathrm{GB}}:=\mathbb{E}_{\Pi_{n}^{\eta}}[\boldsymbol{\theta}]$, and (iii) the one-step Newton update from $\boldsymbol{\theta}_{\mathrm{GB}}$ based on the penalized score. Details are given in Section A.
@@ 3 Main results
We denote by $q_{n}^{\tilde{\boldsymbol{\theta}}_{n}}$ the density of $\sqrt{s_{n}}(\boldsymbol{\theta}-\tilde{\boldsymbol{\theta}}_{n})$ under $\Pi_{n}^{\eta}$, and by $\mathcal{N}(\cdot\mid\boldsymbol{\mu},\boldsymbol{\Sigma})$ the density of the $p$-variate normal distribution with mean $\boldsymbol{\mu}$ and covariance matrix $\boldsymbol{\Sigma}$. Using local asymptotic normality arguments in the spirit of miller2021asymptotic, adapted to our loss-based criterion with effective scale $s_{n}$, to possibly $n$-dependent priors and to non-i.i.d. data, and evaluated at admissible centers, we obtain the following total-variation limit.
@@
$\int_{\mathbb{R}^{p}}\left|q_{n}^{\tilde{\boldsymbol{\theta}}_{n}}(\boldsymbol{x})-\mathcal{N}\!\left.\Bigl(\boldsymbol{x}\right|\boldsymbol{0},\!\left(\eta\boldsymbol{J}_{\lambda}^{\star}\right)^{-1}\Bigr)\right|\mathop{}\!\mathrm{d}\boldsymbol{x}\longrightarrow 0,\quad n\to\infty.$
@@
$\sqrt{s_{n}}\left(\boldsymbol{\theta}_{\mathrm{calib}}^{(d)}-\tilde{\boldsymbol{\theta}}_{n}\right)\longrightarrow_{d}\mathcal{N}\left(\boldsymbol{0},\boldsymbol{V}_{\mathrm{target}}^{\star}\right),\quad n\to\infty.$
@@
Hence the asymptotic law of $\sqrt{s_{n}}(\boldsymbol{\theta}_{\mathrm{calib}}^{(d)}-\tilde{\boldsymbol{\theta}}_{n})$ is invariant to the learning rate $\eta$.
Remark 3.4 (Extension to convex non-smooth penalties).
The main text assumes $\rho\in C^{2}$ near $\boldsymbol{\theta}^{\lambda}$. For convex non-smooth penalties, the results extend by restricting the analysis to an appropriate active set and then applying the same arguments with all matrices restricted accordingly. Details are provided in Section D.
@@ 4 Practical plug-in calibration
$\widehat{\boldsymbol{\Sigma}}_{\mathrm{post}}:=\frac{1}{D}\sum_{d=1}^{D}\bigl(\boldsymbol{\theta}^{(d)}-\hat{\boldsymbol{\theta}}_{\mathrm{GB}}\bigr)\bigl(\boldsymbol{\theta}^{(d)}-\hat{\boldsymbol{\theta}}_{\mathrm{GB}}\bigr)^{\top}.$
@@
$s_{n}\!\left\|\widehat{\boldsymbol{\Sigma}}_{\mathrm{post}}-\boldsymbol{\Sigma}_{\mathrm{post},n}\right\|_{F}\longrightarrow_{p}0,\quad\boldsymbol{\Sigma}_{\mathrm{post},n}:=\operatorname{Var}_{\Pi_{n}^{\eta}}\!\left[\boldsymbol{\theta}\mid\left\{\mathcal{D}_{i}\right\}_{i=1}^{n}\right],$
@@
$s_{n}\widehat{\boldsymbol{\Sigma}}_{\mathrm{post}}\longrightarrow_{p}\boldsymbol{H}_{0}^{-1},\quad n\to\infty,$
@@
$\hat{\boldsymbol{J}}_{\lambda}\to_{p}\boldsymbol{J}_{\lambda}^{\star},\quad\hat{\boldsymbol{V}}_{\mathrm{target}}\to_{p}\boldsymbol{V}_{\mathrm{target}}^{\star}.$
@@
$\hat{\boldsymbol{H}}_{0}^{-1}:=s_{n}\widehat{\boldsymbol{\Sigma}}_{\mathrm{post}},\quad\hat{\boldsymbol{\Omega}}:=\hat{\boldsymbol{V}}_{\mathrm{target}}^{1/2}\hat{\boldsymbol{H}}_{0}^{1/2},\quad\hat{\boldsymbol{\theta}}_{\mathrm{calib}}^{(d)}:=\bar{\boldsymbol{\theta}}_{n}+\hat{\boldsymbol{\Omega}}\bigl(\boldsymbol{\theta}^{(d)}-\hat{\boldsymbol{\theta}}_{\mathrm{GB}}\bigr).$ (4)
@@
$\sqrt{s_{n}}\bigl(\hat{\boldsymbol{\theta}}_{\mathrm{calib}}^{(d)}-\bar{\boldsymbol{\theta}}_{n}\bigr)\longrightarrow_{d}\mathcal{N}\bigl(\boldsymbol{0},\boldsymbol{V}_{\mathrm{target}}^{\star}\bigr),\quad n\to\infty.$
@@ 5.1 Random-intercept linear mixed model with a Huber loss
Let groups $i=1,\dots,G$ have sizes $n_{i}$ and total $n=\sum_{i=1}^{G}n_{i}$. Observations are $(y_{ij},\boldsymbol{x}_{ij})\in\mathbb{R}\times\mathbb{R}^{p}$ with $y_{ij}=\boldsymbol{x}_{ij}^{\top}\boldsymbol{\beta}+b_{i}+\varepsilon_{ij}$, $b_{i}\sim\mathcal{N}(0,\tau^{2})$, $\varepsilon_{ij}\sim\mathcal{N}(0,\sigma^{2})$, independent across $i$. Stack $\boldsymbol{y}_{i}=(y_{i1},\dots,y_{in_{i}})^{\top}$ and $\boldsymbol{X}_{i}\in\mathbb{R}^{n_{i}\times p}$ with rows $\boldsymbol{x}_{ij}^{\top}$. Take the working marginal covariance $\boldsymbol{\Sigma}_{i}=\tau^{2}\boldsymbol{1}_{n_{i}}\boldsymbol{1}_{n_{i}}^{\top}+\sigma^{2}\boldsymbol{I}_{n_{i}}$ and its symmetric square root $\boldsymbol{L}_{i}\boldsymbol{L}_{i}^{\top}=\boldsymbol{\Sigma}_{i}$, and define whitened objects
@@
Let $\Psi(\boldsymbol{\beta})=\mathbb{E}[\boldsymbol{U}_{n}(\boldsymbol{\beta})]$. Let $\boldsymbol{\mu}\in\mathbb{R}^{p}$ be a fixed ridge center and $\boldsymbol{Q}\in\mathbb{R}^{p\times p}$ be a given symmetric positive definite penalty matrix. For the ridge penalty $\rho(\boldsymbol{\beta})=2^{-1}(\boldsymbol{\beta}-\boldsymbol{\mu})^{\top}\boldsymbol{Q}(\boldsymbol{\beta}-\boldsymbol{\mu})$ and $\lambda\in[0,\infty)$, define $\boldsymbol{\beta}^{\lambda}$ by $\boldsymbol{0}=\Psi(\boldsymbol{\beta})+\lambda\nabla\rho(\boldsymbol{\beta})$. Set, at $\boldsymbol{\beta}^{\lambda}$, $\boldsymbol{J}^{\star}=\nabla_{\boldsymbol{\beta}}\Psi(\boldsymbol{\beta}^{\lambda})$, $\boldsymbol{K}^{\star}=\lim_{n\to\infty}n\operatorname{Var}[\boldsymbol{U}_{n}(\boldsymbol{\beta}^{\lambda})]$, $\boldsymbol{J}_{\lambda}^{\star}=\boldsymbol{J}^{\star}+\lambda\boldsymbol{Q}$, and the target sandwich $\boldsymbol{V}_{\mathrm{target}}^{\star}=(\boldsymbol{J}_{\lambda}^{\star})^{-1}\boldsymbol{K}^{\star}(\boldsymbol{J}_{\lambda}^{\star})^{-1}$.
@@
Take a Gaussian prior with $n$-dependent scale $\boldsymbol{\beta}\sim\mathcal{N}(\boldsymbol{\mu},\{(\lambda s_{n})\boldsymbol{Q}\}^{-1})$. The posterior is $\Pi_{n}^{\eta}(\mathop{}\!\mathrm{d}\boldsymbol{\beta})\propto\exp\{-\eta M_{n}(\boldsymbol{\beta})\}\pi_{n}(\boldsymbol{\beta})\mathop{}\!\mathrm{d}\boldsymbol{\beta}$ and the local precision per $s_{n}$ is $\boldsymbol{H}_{0}=\eta\boldsymbol{J}^{\star}_{\lambda}$. With an admissible center $\tilde{\boldsymbol{\beta}}_{n}$, estimate $\hat{\boldsymbol{J}}_{\lambda}=\boldsymbol{J}_{n}(\tilde{\boldsymbol{\beta}}_{n})+\lambda_{n}\boldsymbol{Q}$, $\hat{\boldsymbol{K}}=n^{-1}\sum_{i=1}^{G}\hat{\boldsymbol{U}}_{i}\hat{\boldsymbol{U}}_{i}^{\top}$, $\hat{\boldsymbol{U}}_{i}=-\tilde{\boldsymbol{X}}_{i}^{\top}\psi_{c}(\tilde{\boldsymbol{r}}_{i}(\tilde{\boldsymbol{\beta}}_{n}))$, and $\hat{\boldsymbol{V}}_{\mathrm{target}}=\hat{\boldsymbol{J}}_{\lambda}^{-1}\hat{\boldsymbol{K}}\hat{\boldsymbol{J}}_{\lambda}^{-1}$. From posterior draws $\{\boldsymbol{\beta}^{(d)}\}_{d=1}^{D}$ with mean $\hat{\boldsymbol{\beta}}_{\mathrm{GB}}$, set
@@
$\widehat{\boldsymbol{\Sigma}}_{\mathrm{post}}=\frac{1}{D}\sum_{d=1}^{D}\bigl(\boldsymbol{\beta}^{(d)}-\hat{\boldsymbol{\beta}}_{\mathrm{GB}}\bigr)\bigl(\boldsymbol{\beta}^{(d)}-\hat{\boldsymbol{\beta}}_{\mathrm{GB}}\bigr)^{\top},\quad\hat{\boldsymbol{H}}_{0}^{-1}=n\widehat{\boldsymbol{\Sigma}}_{\mathrm{post}},\quad\hat{\boldsymbol{\Omega}}=\hat{\boldsymbol{V}}_{\mathrm{target}}^{1/2}\hat{\boldsymbol{H}}_{0}^{1/2}.$
@@
Then, calibrated draws are $\boldsymbol{\beta}_{\mathrm{calib}}^{(d)}=\tilde{\boldsymbol{\beta}}_{n}+\hat{\boldsymbol{\Omega}}\bigl(\boldsymbol{\beta}^{(d)}-\hat{\boldsymbol{\beta}}_{\mathrm{GB}}\bigr)$. By Proposition 4.3, $\sqrt{s_{n}}(\boldsymbol{\beta}_{\mathrm{calib}}^{(d)}-\tilde{\boldsymbol{\beta}}_{n})\to_{d}\mathcal{N}(\boldsymbol{0},\boldsymbol{V}_{\mathrm{target}}^{\star})$. Details of the augmentation and Markov chain Monte Carlo are deferred to Section E.

### v3

@@ 1 Introduction
In this paper, we revisit the open-faced sandwich adjustment (shaby2014open-faced) and place it in a general asymptotic framework for generalized Bayes based on loss functions with an effective sample size, allowing sample-size-dependent priors, possibly non-i.i.d. observations, and convex penalties under model misspecification. First, we show that, under sample-size-dependent priors, the generalized posterior admits a normal limit in total variation and we establish a prior–penalty correspondence that identifies the target curvature and variability. Second, building on this justification, we improve the open-faced sandwich adjustment from covariance rescaling to a location–scale calibration that is implementable via plug-in sandwich estimators, requiring only posterior draws and empirical moments, with no bootstrap or learning-rate tuning. Third, we prove that the calibrated draws converge to the target distribution for any learning rate, so that the asymptotic uncertainty quantification is invariant to the choice of learning rate. Our empirical illustration for a random-intercept mixed model with Huber loss and Gaussian prior with sample-size-dependent scale shows that the calibrated intervals track frequentist benchmarks across several orders of magnitude in the learning rate, whereas uncalibrated generalized Bayes posteriors exhibit substantial sensitivity.
@@ 2 Problem setup
Fix $p\in\mathbb{N}$ and let $\Theta\subset\mathbb{R}^{p}$ be an open parameter space. Write $\boldsymbol{\theta}\in\Theta$ for the parameter, and let $\|\cdot\|$ denote the Euclidean norm. Let $M_{n}:\Theta\to(-\infty,\infty]$ be an empirical criterion with effective scale $s_{n}\to\infty$; typical examples include i.i.d. additive losses, for which $s_{n}=n$; $m$-variate $U$-type losses with non-degenerate kernels, where $s_{n}=\binom{n}{m}$; and kernel-smoothed losses with bandwidth $h_{n}>0$, in which case $s_{n}=nh_{n}$ and both $h_{n}\to 0$ and $nh_{n}\to\infty$. Define the rescaled criterion and its first and second derivatives by
@@
We write $R(\boldsymbol{\theta}):=E_{P^{\star}}(\bar{M}_{n}(\boldsymbol{\theta}))$ for the corresponding population criterion, where $P^{\star}$ denotes the true data-generating distribution.
@@
This loss-based representation makes explicit that the learning rate $\eta$ scales the entire loss, in line with bissiri2016general. Throughout the main text we assume that $\rho$ is $C^{2}$ in a neighborhood of the target point introduced below, as stated in Assumption 2.1; the convex non-smooth case is treated in Section D.
Let $\mathcal{D}=\{\mathcal{D}_{i}\}_{i=1}^{n}$ be observations from an unknown distribution $P^{\star}$ and $\Psi(\boldsymbol{\theta}):=E_{P^{\star}}(\psi(\mathcal{D}_{1},\boldsymbol{\theta}))$ denote the population estimating function corresponding to $M_{n}$, for some measurable $\psi$. In the $M$-estimation case $M_{n}(\boldsymbol{\theta})=\sum_{i=1}^{n}m(\mathcal{D}_{i},\boldsymbol{\theta})$, one may take $\psi(\mathcal{D},\boldsymbol{\theta})=\nabla_{\boldsymbol{\theta}}m(\mathcal{D},\boldsymbol{\theta})$. The penalized population equation is
@@
$\boldsymbol{J}^{\star}:=\nabla_{\boldsymbol{\theta}}\Psi\bigl(\boldsymbol{\theta}^{\lambda}\bigr),\quad\boldsymbol{K}^{\star}:=\operatorname{var}_{P^{\star}}\!\left(\psi\bigl(\mathcal{D}_{1},\boldsymbol{\theta}^{\lambda}\bigr)\right),\quad\boldsymbol{H}_{\rho}\bigl(\boldsymbol{\theta}^{\lambda}\bigr):=\nabla^{2}\rho\bigl(\boldsymbol{\theta}^{\lambda}\bigr).$
@@
There exists an open neighborhood $\mathcal{N}$ of $\boldsymbol{\theta}^{\lambda}$ such that: (i) the penalized population equation (1) has the unique solution $\boldsymbol{\theta}^{\lambda}\in\mathcal{N}$; (ii) $R\in C^{3}(\mathcal{N})$ and $\sup_{\boldsymbol{\theta}\in\mathcal{N}}\|\nabla^{3}\bar{M}_{n}(\boldsymbol{\theta})\|=O_{p}(1)$; (iii) $\sup_{\boldsymbol{\theta}\in\mathcal{N}}\|\boldsymbol{J}_{n}(\boldsymbol{\theta})-\boldsymbol{J}^{\star}\|\to_{p}0$ and $\boldsymbol{J}^{\star}$ is nonsingular; and (iv) $s_{n}^{1/2}\{\boldsymbol{U}_{n}(\boldsymbol{\theta}^{\lambda})-\Psi(\boldsymbol{\theta}^{\lambda})\}\to_{d}N(\boldsymbol{0},\boldsymbol{K}^{\star})$.
@@
Let $\mathcal{U}_{n}=\{\boldsymbol{\theta}:\|\boldsymbol{\theta}-\boldsymbol{\theta}^{\lambda}\|\leq cs_{n}^{-1/2}\}$ for some fixed constant $c>0$. The prior remainder $r_{n}$ is locally Lipschitz on $\mathcal{U}_{n}$ with Lipschitz constant $L_{n}=o_{p}(s_{n}^{1/2})$.
@@
Under Assumptions 2.12.4, the following centers are admissible: (i) any maximum a posteriori (MAP) estimator, (ii) the posterior mean $\boldsymbol{\theta}_{\mathrm{GB}}:=E_{\Pi_{n}^{\eta}}(\boldsymbol{\theta})$, and (iii) the one-step Newton update from $\boldsymbol{\theta}_{\mathrm{GB}}$ based on the penalized score. Details are given in Section A.
@@ 3 Main results
We denote by $q_{n}^{\tilde{\boldsymbol{\theta}}_{n}}$ the density of $s_{n}^{1/2}(\boldsymbol{\theta}-\tilde{\boldsymbol{\theta}}_{n})$ under $\Pi_{n}^{\eta}$, and by $N(\cdot\mid\boldsymbol{\mu},\boldsymbol{\Sigma})$ the density of the $p$-variate normal distribution with mean $\boldsymbol{\mu}$ and covariance matrix $\boldsymbol{\Sigma}$. Using local asymptotic normality arguments in the spirit of miller2021asymptotic, adapted to our loss-based criterion with effective scale $s_{n}$, to possibly $n$-dependent priors and to non-i.i.d. data, and evaluated at admissible centers, we obtain the following total-variation limit.
@@
$\int_{\mathbb{R}^{p}}\left|q_{n}^{\tilde{\boldsymbol{\theta}}_{n}}(\boldsymbol{x})-N\!\left.\Bigl(\boldsymbol{x}\right|\boldsymbol{0},\!\left(\eta\boldsymbol{J}_{\lambda}^{\star}\right)^{-1}\Bigr)\right|\mathop{}\!\mathrm{d}\boldsymbol{x}\longrightarrow_{p}0,\quad n\to\infty.$
@@
$s_{n}^{1/2}\left(\boldsymbol{\theta}_{\mathrm{calib}}^{(d)}-\tilde{\boldsymbol{\theta}}_{n}\right)\longrightarrow_{d}N\!\left(\boldsymbol{0},\boldsymbol{V}_{\mathrm{target}}^{\star}\right),\quad n\to\infty.$
@@
Hence the asymptotic law of $s_{n}^{1/2}(\boldsymbol{\theta}_{\mathrm{calib}}^{(d)}-\tilde{\boldsymbol{\theta}}_{n})$ is invariant to the learning rate $\eta$.
@@ 4 Practical plug-in calibration
$\hat{\boldsymbol{\Sigma}}_{\mathrm{post}}:=\frac{1}{D}\sum_{d=1}^{D}\bigl(\boldsymbol{\theta}^{(d)}-\hat{\boldsymbol{\theta}}_{\mathrm{GB}}\bigr)\bigl(\boldsymbol{\theta}^{(d)}-\hat{\boldsymbol{\theta}}_{\mathrm{GB}}\bigr)^{\top}.$
@@
$s_{n}\!\left\|\hat{\boldsymbol{\Sigma}}_{\mathrm{post}}-\boldsymbol{\Sigma}_{\mathrm{post},n}\right\|_{F}\longrightarrow_{p}0,\quad\boldsymbol{\Sigma}_{\mathrm{post},n}:=\operatorname{var}_{\Pi_{n}^{\eta}}\!\left(\boldsymbol{\theta}\mid\mathcal{D}\right),$
@@
$s_{n}\hat{\boldsymbol{\Sigma}}_{\mathrm{post}}\longrightarrow_{p}\boldsymbol{H}_{0}^{-1},\quad n\to\infty,$
@@
$\hat{\boldsymbol{J}}_{\lambda}\longrightarrow_{p}\boldsymbol{J}_{\lambda}^{\star},\quad\hat{\boldsymbol{V}}_{\mathrm{target}}\longrightarrow_{p}\boldsymbol{V}_{\mathrm{target}}^{\star}.$
@@
$\hat{\boldsymbol{H}}_{0}^{-1}:=s_{n}\hat{\boldsymbol{\Sigma}}_{\mathrm{post}},\quad\hat{\boldsymbol{\Omega}}:=\hat{\boldsymbol{V}}_{\mathrm{target}}^{1/2}\hat{\boldsymbol{H}}_{0}^{1/2},\quad\hat{\boldsymbol{\theta}}_{\mathrm{calib}}^{(d)}:=\bar{\boldsymbol{\theta}}_{n}+\hat{\boldsymbol{\Omega}}\bigl(\boldsymbol{\theta}^{(d)}-\hat{\boldsymbol{\theta}}_{\mathrm{GB}}\bigr).$ (4)
@@
$s_{n}^{1/2}\bigl(\hat{\boldsymbol{\theta}}_{\mathrm{calib}}^{(d)}-\bar{\boldsymbol{\theta}}_{n}\bigr)\longrightarrow_{d}N\bigl(\boldsymbol{0},\boldsymbol{V}_{\mathrm{target}}^{\star}\bigr),\quad n\to\infty.$
@@ 5.1 Random-intercept linear mixed model with a Huber loss
Let groups $i=1,\ldots,G$ have sizes $n_{i}$ and total $n=\sum_{i=1}^{G}n_{i}$. Observations are $(y_{ij},\boldsymbol{x}_{ij})\in\mathbb{R}\times\mathbb{R}^{p}$ with $y_{ij}=\boldsymbol{x}_{ij}^{\top}\boldsymbol{\beta}+b_{i}+\varepsilon_{ij}$, $b_{i}\sim N(0,\tau^{2})$, $\varepsilon_{ij}\sim N(0,\sigma^{2})$, independent across $i$. Stack $\boldsymbol{y}_{i}=(y_{i1},\ldots,y_{in_{i}})^{\top}$ and $\boldsymbol{X}_{i}\in\mathbb{R}^{n_{i}\times p}$ with rows $\boldsymbol{x}_{ij}^{\top}$. Take the working marginal covariance $\boldsymbol{\Sigma}_{i}=\tau^{2}\boldsymbol{1}_{n_{i}}\boldsymbol{1}_{n_{i}}^{\top}+\sigma^{2}\boldsymbol{I}_{n_{i}}$ and its symmetric square root $\boldsymbol{L}_{i}\boldsymbol{L}_{i}^{\top}=\boldsymbol{\Sigma}_{i}$, and define whitened objects
@@
Let $\Psi(\boldsymbol{\beta})=E(\boldsymbol{U}_{n}(\boldsymbol{\beta}))$. Let $\boldsymbol{\mu}\in\mathbb{R}^{p}$ be a fixed ridge center and $\boldsymbol{Q}\in\mathbb{R}^{p\times p}$ be a given symmetric positive definite penalty matrix. For the ridge penalty $\rho(\boldsymbol{\beta})=2^{-1}(\boldsymbol{\beta}-\boldsymbol{\mu})^{\top}\boldsymbol{Q}(\boldsymbol{\beta}-\boldsymbol{\mu})$ and $\lambda\in[0,\infty)$, define $\boldsymbol{\beta}^{\lambda}$ by $\boldsymbol{0}=\Psi(\boldsymbol{\beta})+\lambda\nabla\rho(\boldsymbol{\beta})$. Set, at $\boldsymbol{\beta}^{\lambda}$, $\boldsymbol{J}^{\star}=\nabla_{\boldsymbol{\beta}}\Psi(\boldsymbol{\beta}^{\lambda})$, $\boldsymbol{K}^{\star}=\lim_{n\to\infty}n\operatorname{var}(\boldsymbol{U}_{n}(\boldsymbol{\beta}^{\lambda}))$, $\boldsymbol{J}_{\lambda}^{\star}=\boldsymbol{J}^{\star}+\lambda\boldsymbol{Q}$, and the target sandwich $\boldsymbol{V}_{\mathrm{target}}^{\star}=(\boldsymbol{J}_{\lambda}^{\star})^{-1}\boldsymbol{K}^{\star}(\boldsymbol{J}_{\lambda}^{\star})^{-1}$.
@@
Take a Gaussian prior with $n$-dependent scale $\boldsymbol{\beta}\sim N(\boldsymbol{\mu},\{(\lambda s_{n})\boldsymbol{Q}\}^{-1})$. The posterior is $\Pi_{n}^{\eta}(\mathop{}\!\mathrm{d}\boldsymbol{\beta})\propto\exp\{-\eta M_{n}(\boldsymbol{\beta})\}\pi_{n}(\boldsymbol{\beta})\mathop{}\!\mathrm{d}\boldsymbol{\beta}$ and the local precision per $s_{n}$ is $\boldsymbol{H}_{0}=\eta\boldsymbol{J}^{\star}_{\lambda}$. With an admissible center $\tilde{\boldsymbol{\beta}}_{n}$, estimate $\hat{\boldsymbol{J}}_{\lambda}=\boldsymbol{J}_{n}(\tilde{\boldsymbol{\beta}}_{n})+\lambda_{n}\boldsymbol{Q}$, $\hat{\boldsymbol{K}}=n^{-1}\sum_{i=1}^{G}\hat{\boldsymbol{U}}_{i}\hat{\boldsymbol{U}}_{i}^{\top}$, $\hat{\boldsymbol{U}}_{i}=-\tilde{\boldsymbol{X}}_{i}^{\top}\psi_{c}(\tilde{\boldsymbol{r}}_{i}(\tilde{\boldsymbol{\beta}}_{n}))$, and $\hat{\boldsymbol{V}}_{\mathrm{target}}=\hat{\boldsymbol{J}}_{\lambda}^{-1}\hat{\boldsymbol{K}}\hat{\boldsymbol{J}}_{\lambda}^{-1}$. From posterior draws $\{\boldsymbol{\beta}^{(d)}\}_{d=1}^{D}$ with mean $\hat{\boldsymbol{\beta}}_{\mathrm{GB}}$, set
@@
$\hat{\boldsymbol{\Sigma}}_{\mathrm{post}}=\frac{1}{D}\sum_{d=1}^{D}\bigl(\boldsymbol{\beta}^{(d)}-\hat{\boldsymbol{\beta}}_{\mathrm{GB}}\bigr)\bigl(\boldsymbol{\beta}^{(d)}-\hat{\boldsymbol{\beta}}_{\mathrm{GB}}\bigr)^{\top},\quad\hat{\boldsymbol{H}}_{0}^{-1}=n\hat{\boldsymbol{\Sigma}}_{\mathrm{post}},\quad\hat{\boldsymbol{\Omega}}=\hat{\boldsymbol{V}}_{\mathrm{target}}^{1/2}\hat{\boldsymbol{H}}_{0}^{1/2}.$
@@
Then, calibrated draws are $\boldsymbol{\beta}_{\mathrm{calib}}^{(d)}=\tilde{\boldsymbol{\beta}}_{n}+\hat{\boldsymbol{\Omega}}\bigl(\boldsymbol{\beta}^{(d)}-\hat{\boldsymbol{\beta}}_{\mathrm{GB}}\bigr)$. By Proposition 4.3, $s_{n}^{1/2}(\boldsymbol{\beta}_{\mathrm{calib}}^{(d)}-\tilde{\boldsymbol{\beta}}_{n})\to_{d}N(\boldsymbol{0},\boldsymbol{V}_{\mathrm{target}}^{\star})$. Details of the augmentation and Markov chain Monte Carlo are deferred to Section E.